OpenStudy (macaylam):

The center of an ice rink is located at (0, 0) on a coordinate system measured in feet. Sandi skates along a path that can be modeled by the equation y = 0.08x2 - 1.6x + 13. David starts at (20, 13) and skates along a path that can be modeled by a quadratic function with a vertex at (10, 5). Which system of equations models the paths of the two skaters? y = 0.08x2-1.6x+13 and y = 0.08(x-10)2+5 y = 0.08x2-1.6x+13 and y = 1.6(x-10)2+5 y = 0.08x2-1.6x+13 and y = 0.8(x-10)2+5

OpenStudy (triciaal):

can you write the second equation ? y = a(x-h)^2 + k vertex at (h,k) h = -b/2a and k = f(h) passes through (20, 13)